OpenStudy (anonymous):

The face of a 12 sided die are numbered from 1 to 12. What is the probability of rolling 9 AT LEAST TWICE in ten tries. Medal for answer

OpenStudy (atlas):

Can you find out how many ways the thrower will get 9 twice or more than twice times?

OpenStudy (atlas):

Hint : the user can get 9 exactly twice in 12 throws by 12C2

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

no, it only says 9 at least twice in 10 tries

OpenStudy (anonymous):

doesnt say how many ways it can b thrown

OpenStudy (the_fizicx99):

I think it could be favorable outcomes / number of outcomes

OpenStudy (anonymous):

its kind of like that, its a success/ failure probability

OpenStudy (the_fizicx99):

Plug in your #'s?

OpenStudy (anonymous):

its more complicated than that

OpenStudy (anonymous):

what is the probability of getting 0 nines

OpenStudy (the_fizicx99):

Do you have options to chose from?

OpenStudy (anonymous):

no, but u can use binomial distribution to help you

OpenStudy (anonymous):

P(at least 2) = 1- [P(0) + P(1)] P is probability

OpenStudy (anonymous):

is P(0)= 11/12?

OpenStudy (anonymous):

i thought of that method too, but i didnt what probability of 0 nine is

OpenStudy (anonymous):

you can find P(0) and P(1) with binomial distribution (nCk)*(p^k)*(q^(n-k)) p is probability of succses = 1/12 q is probability of failure = 9/12

OpenStudy (anonymous):

oops sory q = 11/12

OpenStudy (the_fizicx99):

Same thing as Success/Failure 9/12 and 11/12

OpenStudy (anonymous):

I got zero for the answer, but its wrong

OpenStudy (the_fizicx99):

Did you get a decimal?

OpenStudy (anonymous):

\[p(getting 9 inone throw)=\frac{ 1 }{ 12 }\] \[P \prime=1-\frac{ 1 }{ 12 }=\frac{ 11 }{12 }\] \[p(no 9 )=\left( \frac{ 11 }{ 12} \right)^{10}\]\[P \left( one 9 only \right)=\frac{ 1 }{12 }\left( \frac{ 11 }{ 12 } \right)^{9}\]

OpenStudy (anonymous):

Required probability=1-[P(no 9)+P(one 9 only)]

OpenStudy (anonymous):

Im not getting the right answer, what are you getting?

OpenStudy (anonymous):

the answer should b 0.2003

OpenStudy (anonymous):

what's your answer? P(0) = (10C0)*((1/12)^0)*(11/12)^10 P(1) = (10C1)*((1/12)^1)*(11/12)^9

OpenStudy (anonymous):

there we go :DD

OpenStudy (anonymous):

but i dont get how you did that

OpenStudy (anonymous):

can you explain it plz

OpenStudy (anonymous):

C is combination and for calculate (11/12)^9 and (11/12)^10 i use calculator :D

OpenStudy (anonymous):

why did u put the exponent of 9 and 10

OpenStudy (mathmale):

For those who own a TI-83 or TI-84 calculator and are familiar with probability distributions: type in the following: binomcdf(10,1/2,1). Result: 0.7997; subtracting this from 1 results in 0.2003, which is, as you say, the correct answer.

OpenStudy (mathmale):

For those who don't own one of these calculators, or who must "show work," use the formula for binomial probability:

OpenStudy (anonymous):

n is number of trials and k is number of possible ways to choose k "successes" from n trials so n = 10 and k=0 for P(0) , k=1 for P(1)

OpenStudy (anonymous):

P(k) = (nCk)*(p^k)*(q^(n-k))

OpenStudy (mathmale):

hanifahrika is right on target, calculating, as she did, the probabilities of getting zero 9's in 10 trials and one 9 in 10 trials. Add these two probabilities together to get the probability of getting either zero or one 9's. To get the probability of getting at least 2 9's, subtract this sum from 1: 1-0.7997=0.2003.

OpenStudy (anonymous):

is that the same formula as P(x)= nCr P^r q^n-r?

OpenStudy (mathmale):

Yes, it is. I'd suggest typing p(x) = nCr * (p^r) * (q^[n-r]) for the sake of clarity.

OpenStudy (anonymous):

Okay ty guys I understand it now

OpenStudy (mathmale):

That's essentially the same as what hanifahrika provided. Delighted!